OpenStudy (anonymous):

Help pleaaaseeee.. 1) Find the balance in the account $2000 principle earning 7% compounded semi-annually after 25 years. 2)A tractor costs $14,340 and depreciates in value by 15% per year. How much will the tractor be worth after 3 years?

OpenStudy (anonymous):

For a) \[A=P(1+r/n)^{nt} \] where P is the principal investment, n is the number of times your principal is compounded annually , and r is the percentage increase as a decimal. So we have P=2000 , n=2, t=25, and r=.07

OpenStudy (anonymous):

so we're gonna write it like this 2000(1+.07/2)^25 ? @shawmoes

OpenStudy (kropot72):

@proud_yemeniah The expression in brackets must be raised to the power of nt where n = 2 and t = 25. In your substitution you have not considered n when writing the value of the power.

OpenStudy (anonymous):

i don't really understand what im suppose to do @kropot72

OpenStudy (kropot72):

Substituting into the formula given by @shawmoes gives: \[\large A=2000(1+\frac{0.07}{2})^{(2\times25)}\]

OpenStudy (anonymous):

we then use the PEDMAS rule right ? @kropot72

Nnesha (nnesha):

first solve parentheses which is exponent

Nnesha (nnesha):

are you there @proud_yemeniah_

OpenStudy (anonymous):

yes, so its gonna be 2000(1+0.07/2)^50

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

so then is it gonna be like 2000*1 .07/2^25 ?

OpenStudy (anonymous):

@shawmoes

Nnesha (nnesha):

nope you can't just add 1 and .07 bec \[(1+ \frac{ .07 }{ 2 })\] first you have to find common denominator

OpenStudy (anonymous):

i don't really understand @Nnesha

Nnesha (nnesha):

 multiply 2 by 1 and then add .07